2. At a certain temperature a saturated solution of BaF2 has a concentration of 4.59 x 10-2 M. What is the Ksp for BaF2 at this temperature? (3 marks)



3. What is the concentration of the ions present in a saturated Cd(OH)2 solution? Ksp = 5.3 x 10-15. (4 marks)


1
Expert's answer
2022-05-06T14:51:02-0400

2.BaF<===> Ba2+ + 2F-

Ksp = [Ba2+][2F-]2

Ksp = 4.59 x 10-2 M * (2•4.59 x 10-2 M)2 = 3.87x10-4

3.Cd(OH)2 <===> Cd2+ + 2OH-

Ksp = [Cd+][2OH-]2

[Cd+] - x; [OH-]2 - 2x

5.3 x 10-15 = x • (2x)2

5.3 x 10-15 = 4x3

x3 = 1.325 x 10-15

x = 1.098 x 105

[Cd+] - 1.098 x 105M

[OH-]2 - (2•1.098 x 105) = 2.196x105 M